# Ego vs. Exo and Active vs. Passive: Investigating the Effects of Viewpoint and Navigation on Spatial Immersion and Understanding in Immersive Storytelling

### v1

@@ 4.1. Design Consideration Exploration
DC3 (navigation trigger) in passive designs was triggered by simple actions like button presses. In active designs, it involved either locomotions (Di Luca et al., 2021) (e.g., walking or teleporting) or physical interactions (e.g., opening a window in VR).
@@ 4.2. Case 1: How 911 changed Financial Distract in Manhattan
A VR adaptation of a story about lower Manhattan’s transformation post-9/11. It displays four screenshots from a VR environment where the story is adapted with interactive 3D elements. In the VR environment, participants engage with the following components: N1 (Main Visual Element, showing lower Manhattan in 3D), N2 (Story Narratives and Text Guidance displayed on a floating panel), N3 (Active Navigation, allowing users to manipulate the environment), and N4 (Historical imagery overlayed on the 3D map). The VR version integrates interactive spatial elements and narrative text for a more immersive experience.
This story covers the societal changes in Manhattan’s Financial District post-9/11 (Bentley, 2021). The original story featured a 3D model of Manhattan’s densely packed buildings, making it ideal for an exocentric VR adaptation. The story has 15 steps and 611 words in total. A step includes a paragraph of story narratives with corresponding visual elements. Audiences procceed to next step by following a specific navigation trigger.
Main Visual Element. We built a 3D landscape of Manhattan in VR (Figure 5-N1) through Mapbox (Mapbox, 2017). The landscape is 2m-by-2m square base map by default. The current building information came from mapbox.3d-buildings tileset. When the story narratives mentioned a particular group of buildings (e.g. WTC buildings), they are highlighted in blue.
@@ 4.3. Case 2: In the Atlantic Ocean, Subtle Shifts Hint at Dramatic Dangers
A VR adaptation of a story about the weakening Gulf Stream in the Atlantic Ocean. It displays four VR screenshots: A1 (Main Visual Element, showing the Earth with a visual representation of the Gulf Stream and story narratives/text guidance), A2 (Active Navigation, where users can manipulate the globe for a better view), A3 (Visual Guidance, highlighting a specific region of the Gulf Stream), and A4 (Text and Visual Guidance, where users can interact with the 3D visualization to explore the story). The VR adaptation offers an interactive, spatial representation of the Gulf Stream data for a more immersive understanding.
This story explores the impact of climate change on the Gulf Stream (Velasquez-Manoff and White, 2021). We only selected the first half of the story, which has the most concentrated 3D spatial information. The main visualization of the original story was a globe with animated current circulations in Atlantic Ocean. Thus, we implemented a corresponding exocentric VR story. We summarized the text-only portion (484 words) into 2 paragraphs (100 words) to keep the length consistency. The story has 15 steps and 503 words in total.
@@ 4.4. Case 3: What the Tulsa Race Massacre Destroyed
A VR adaptation of the story of the Tulsa Race Massacre. It features four VR screenshots: T1 (Text Guidance showing a narrative about the event), T2 (A historical photograph presented in the VR environment), T3 (Main Visual Element with active navigation using teleportation to explore the reconstructed neighborhood), and T4 (Story Narratives displayed as floating panels alongside the VR scene). The VR adaptation allows users to actively explore and interact with visual elements of the historical scene for an immersive learning experience.
The story of this case illustrates the business achievement of black Americans in Tulsa, Oklahoma in the early 1900s and how the town was destroyed by racist mobs (Parshina-Kottas et al., 2021). We only selected the first half the story, which has the most concentrated 3D information. We summarized the second half of the story into a paragraph (130 words) and put it at the end so that story narratives are logically complete while keeping similar length with previous cases. The main visualization of the original story comprises a 3D reconstruction of Tulsa before the massacre, which leads to our design of a VR story with an egocentric viewpoint in a main avenue (Greenwood Ave) of Tulsa (Figure 7-T3). The story has 18 steps and 618 words in total.
@@ 4.5. Case 4: Why Opening Windows Is a Key to Reopening Schools
A VR adaptation of a story on the importance of opening windows for safe school reopening. It features four VR screenshots: C1 (Main Visual Element showing the virtual classroom with a seated arrangement of students), C2 (Active Navigation via teleportation with visual guidance lines showing airflow), C3 (Story Narratives and Text Guidance displayed alongside airflow visualizations), and C4 (Active Navigation through object manipulation, with users grabbing and interacting with windows to alter the airflow in the classroom). The VR version allows for an interactive, spatial understanding of how air circulates within the environment and the impact of opening windows.
@@ 5.1. Study Factors and Conditions
A comparison table outlining the characteristics of egocentric (Ego) vs. exocentric (Exo) viewpoints and active vs. passive navigation in immersive storytelling. On the left, the table compares Ego and Exo based on four characteristics: surrounding visual elements (Ego includes surrounding elements while Exo does not), scale of story scenes (Ego has larger scenes), spatial information density (higher in Exo), and out-of-view target visibility (available in Ego but not Exo). On the right, the table compares Active and Passive navigation in terms of required embodied interactions (more in Active), precise interaction (more in Active), level of control (higher in Active), and required visual search (more in Active than Passive). Larger circles indicate greater intensity or significance of the characteristic.
@@ 5.5. Task Data Collection
Through the study, we aimed to investigate the impact of ego/exocentric viewpoints and active/passive navigation on spatial immersion and understanding of VR stories. For spatial immersion, we collected responses to the Presence Questionnaire (Berkman and Çatak, 2021) and NASA-TLX (Hart and Staveland, 1988) to measure spatial presence and perceived workload, respectively.
@@ 5.6. Hypotheses
Story Session Duration. We hypothesize that active navigation will result in longer story session durations compared to passive navigation. The interactive elements in active navigation require users to engage physically with the environment, adding time to the experience. Prior studies have shown that increased interaction in VR generally extends session time due to higher engagement levels and the need for precise actions in immersive environments (Lages and Bowman, 2018; Servotte et al., 2020). This additional engagement likely leads participants to spend more time exploring the environment and interacting with the content, especially in active conditions.
Additionally, Ego viewpoints are likely to further increase session duration, as spatial navigation in egocentric environments tends to be more complex and challenging, requiring users to explore the scene more actively, which can lead to extended engagement (Van Loon et al., 2018).
Story Understanding. We hypothesize that active navigation will outperform passive navigation on both content comprehension and spatial information depiction, as embodied interaction has been shown to improve cognitive engagement and memory retention (Wilson, 2002; Makransky et al., 2019). Active navigation encourages physical movement, which enhances attention to both content and spatial information, leading to better comprehension and recall (Servotte et al., 2020).
We also expect that exocentric viewpoints will result in better content comprehension because they provide an overview of the entire scene, facilitating the integration of key story elements and improving factual recall (Yang et al., 2018). Egocentric viewpoints will lead to better spatial understanding, as being immersed in the scene helps participants build stronger spatial awareness and memory, which is supported by the ”memory palace” technique (Legge et al., 2012; Krokos et al., 2019).
Spatial Immersion. We expect that Ego viewpoints combined with active navigation will result in the highest levels of spatial immersion. Active navigation increases the sense of presence and control, while Ego viewpoints create a fully immersive experience by surrounding participants with 3D elements (Slater, 2009). Prior studies suggest that immersion is enhanced when users feel directly involved in the environment, with Ego viewpoints offering a more embodied experience (Cummings and Bailenson, 2016).
While active navigation may increase perceived workload due to the effort required to interact with the environment, we hypothesize that the enhanced sense of presence and engagement outweighs the additional cognitive demand, especially when paired with the Ego viewpoint (Hart and Staveland, 1988). This combination offers an optimal balance between immersion and interactivity, despite a potential trade-off with workload (Servotte et al., 2020).
@@ 6.1. Story Session Durations
A bar chart comparing the completion time (in seconds) for different combinations of viewpoint (Egocentric vs. Exocentric) and navigation style (Active vs. Passive) in a VR storytelling study. The first chart shows similar completion times for Ego (blue) and Exo (orange) viewpoints, averaging around 450 seconds. The second chart shows a significant difference between Active (dark gray) and Passive (light gray) navigation, with Active taking longer ( 550 seconds) compared to Passive ( 400 seconds), as indicated by three asterisks *** denoting high statistical significance. The third chart breaks down completion times by both viewpoint and navigation combinations, with Ego Active taking the longest ( 550 seconds), followed by Exo Active, while Passive conditions (Ego Passive and Exo Passive) take less time ( 400 seconds). Significant differences between conditions are marked with two ** or three asterisks *** indicating different levels of statistical significance.
@@ 6.2. Story Understandings
A series of bar charts displaying scores for Main Idea Memorization, Detail Memorization, and Random Inspirations in relation to egocentric (Ego) vs. exocentric (Exo) viewpoints and active vs. passive navigation in a VR storytelling study. For Main Idea Memorization, the first row of charts shows no significant difference between Ego and Exo, or Active and Passive, with both averaging similar scores. When broken down by viewpoint and navigation, Exo Passive has the highest score, while Ego Active and Exo Active also perform well. For Detail Memorization, there is no marked difference between Ego and Exo or Active and Passive, with all combinations performing similarly. For Random Inspirations, the charts show slightly higher scores for Ego over Exo, and Active over Passive. Ego Active has the highest score, suggesting this combination encourages more creative inspiration. Error bars indicate standard deviation across the different conditions.
@@
A series of bar charts showing scores for Object Existence and Object Transform tasks in a VR storytelling study comparing egocentric (Ego) vs. exocentric (Exo) viewpoints and active vs. passive navigation. For Object Existence, the first set of charts shows no significant difference between Ego and Exo or Active and Passive, with all conditions scoring similarly around 2-3 points. In the breakdown by viewpoint and navigation, Exo Passive has the highest score, but differences remain minimal. For Object Transform, the second set of charts reveals a significant difference between Ego and Exo, with Ego scoring higher, particularly in Ego Active, which has the highest score. A single asterisk * indicates statistical significance in the difference between these conditions. Error bars represent standard deviations for each condition.
@@ 6.3. Spatial Immersion
A series of bar charts illustrating scores for Realism, Possibility to Act, Quality of Interface, Possibility to Examine, and Self-Evaluation of Performance in a VR storytelling study, comparing egocentric (Ego) vs. exocentric (Exo) viewpoints and active vs. passive navigation. For Realism, there is a significant difference in Active vs. Passive conditions, with Passive scoring higher as indicated by two asterisks **, while Exo Active scored significantly lower than Exo Passive as marked by one asterisk *. For Possibility to Act, Active outperforms Passive with three asterisks ***, and Ego Active scores the highest among all combinations. Quality of Interface shows little difference between Ego and Exo, or Active and Passive, across all conditions. For Possibility to Examine, Passive scores slightly higher than Active with one asterisk *, and Ego Active scores highest overall. Self-Evaluation of Performance reveals significant differences, with Ego Active performing better than Exo Active and Passive conditions, indicated by one and three asterisks * and ***. Error bars represent standard deviations across conditions.
@@
A series of bar charts displaying scores for Mental Demand, Physical Demand, Temporal Demand, Performance, Effort, and Frustration in a VR storytelling study, comparing egocentric (Ego) vs. exocentric (Exo) viewpoints and active vs. passive navigation. For Mental Demand, there is a significant difference between Active and Passive, with Passive scoring higher, indicated by two asterisks **, while Exo Active shows a significantly higher mental demand as marked by one and two asterisks * and **. Physical Demand reveals a substantial difference between Active and Passive, with Active scoring much higher, as shown by two and three asterisks ** and ***, particularly in the Ego Active condition. Temporal Demand shows a slight difference, with Passive scoring higher than Active, marked by one asterisk *. For Performance, Passive scores slightly higher than Active, with Exo Passive performing the best, marked by one asterisk *. Effort shows higher scores for Exo Active, with significant differences between conditions, marked by three asterisks ***. Finally, Frustration is significantly higher in Exo Active and Passive conditions, marked by one and three asterisks * and ***. Error bars represent standard deviations across the different conditions.
@@ 7.1. How does viewpoints influence story memorability?
Our study shows that the viewpoint is a more decisive factor than the navigation in story memorability. The exocentric viewpoint is slightly better than the egocentric one in the overview and detail memorization. This result matches the disccovery of some prior works, in which the exocentric viewpoint benefits analytical tasks in immersive environments (Yang et al., 2018; Kraus et al., 2019; Yang et al., 2020a), but the advantage in our study is not as significant. Our anticipated reason is that the exocentric viewpoint provide comprehensive information within audiences’ view, which alleviates the problem of targeting story elements in space. However, this also causes a larger information density and more occlusions than the egocentric viewpoint. The relatively smaller scale of visual representations in exo viewpoints increases the attention to the story textual content received and prolongs the retention time in the memory so that they can did a better content comprehension. However, at the same time, audiences could become less sensitive to content transitions and therefore they might miss some important details. Overall, our results indicate that the exocentric viewpoint can be a better design if the primary goal is to promote the understanding and retention of narratives, and aforementioned problems should be considered to further improve the story.
In contrast, the egocentric viewpoint demonstrated better memorability on spatial information than exocentric viewpoint. Similar result is also found in Krokos et al.’s (Krokos et al., 2019) and Yang et al.’s (Yang et al., 2020b) work. Objects in an egocentric viewpoin are often larger and more visible than the exocentric one. Resultedly, audiences can better recognize transforms (i.e. position, rotation and scale) and capture their changes. More importantly, audiences are involved as part of a egocentric story. This enable them to use themselves as a frame of reference to infer transforms. In exocentric stories, they have to refer to one central element (e.g. the base map in subsection 4.2) whose signficant transform change can greatly distract the spatial information perception since audiences need to find a new frame of reference every time. Thus, the egocentric viewpoint is beneficial for those authors who want to convey understandings and perceptions related to the changes 3D spatial transforms (e.g. the renovation process of an office).
@@ 7.2. How does navigation influence spatial immersion?
This finding aligns with Lages et al.’s study (Lages and Bowman, 2018), which showed that walking-based navigation improved performance for users who lacked proficiency in VR manipulation. Our results suggest that the need for higher precision in finding specific angles in exocentric scenarios (e.g., subsection 4.2, subsection 4.3, and the grab-and-move tasks in Lages et al.’s study) may explain the increased workload. The “changing frame of reference” issue discussed in subsection 7.1 adds to the demand for spatial awareness and motor skills to achieve the correct angle, resulting in higher effort and frustration, as reflected in the study.
@@
When designing navigation for VR stories, authors should consider the story’s viewpoint based on these findings. However, focusing solely on individual navigation steps is insufficient. Ferguson et al. (Ferguson et al., 2020), in their investigation of VR educational games, recommend that active navigation should seamlessly integrate with the story content and involve minimal learning curves. We argue that natural active navigation in VR storytelling should combine story progression with user exploration.
@@ 8. Limitations and Future Work
We adapted four selected web-based stories into immersive environments. To guide this adaptation, we conducted a formative user study to identify critical design considerations and explored various design options. Our exploration was not exhaustive; it aimed to facilitate the adaptation process by focusing on standard, off-the-shelf VR solutions. This project concentrated on specific aspects of viewpoint and navigation: ego vs. exo and active vs. passive. While we believe these are unique and essential areas to investigate, we acknowledge a much broader design space for immersive storytelling beyond these aspects. For instance, more sophisticated VR locomotion techniques (Di Luca et al., 2021), such as redirected walking, could enhance the active navigation experience. Similarly, advanced visual guidance methods (Lange et al., 2020), such as those for out-of-view scenarios (Petford et al., 2019; Lin et al., 2021) and multisensory approaches (Melo et al., 2020), offer additional potential for enhancing user experience. We did not incorporate these techniques, as our primary focus was the user study and providing easily accessible techniques available on standard platforms. Building a comprehensive design space to inform future immersive story design is crucial, and we consider this an area for future work.
@@
Our study materials were adapted from web-based stories that were initially designed for web browsers and contained design patterns specifically tailored for web environments (Bach et al., 2018). It is unclear whether these design patterns are still applicable in immersive environments or whether we need to develop unique and more appropriate design patterns for immersive stories. Studying the adaptation or innovation of these design patterns for immersive stories has the potential to bring about a more native immersive experience. Furthermore, the web-based stories we adapted primarily follow a linear, single-scene structure. However, complex stories with multi-path or branching narratives also exist. While our studied scenario can serve as a foundational unit for these more complex stories, and our findings should be partially applicable, especially within their individual scenes. There is a need to explore transitions between scenes and more intricate active navigation, such as manipulating and selecting narrative branches.
@@
Through our study, we found that exocentric and egocentric viewpoints each have their own strengths and limitations. We considered viewpoint as an exclusive design element, testing it in a controlled manner. There is a great opportunity to design immersive stories that leverage both viewpoints to complement each other. In movies and data videos, cinematic techniques like camera movements and zooms have been used to switch focuses and direct audiences’ attention (Amini et al., 2015; Xu et al., 2022; Li et al., 2023). Similar incoporation such as viewing some story pieces from an exocentric view and others from an egocentric view is also promising to investigate. Our study results can provide guidelines on which parts of the story to use which viewpoint. To achieve this, smooth transition techniques are necessary to allow seamless and meaningful switching between viewpoints, which we aim to explore in our next research phase.
@@
Finally, one important aspect of active vs. passive interaction is the agency of the readers. In our study, readers were given the freedom to navigate by themselves with guidance. However, there are other types of freedom that could be provided to readers, such as allowing them to freely decide their focus or even rearrange the storyline. This may empower readers but also risks them getting lost, particularly due to the more flexible interactivity in immersive environments. Balancing exploration and narrative control is a fundamental question in storytelling (Thudt et al., 2018). Studying the appropriate threshold or mechanism for this balance is an important research thread in immersive storytelling.

### v2

@@ 4.1. Design Consideration Exploration
DC3 (navigation trigger) in passive designs was triggered by simple actions like button presses. In active designs, it involved either locomotions (di2021locomotion) (e.g., walking or teleporting) or physical interactions (e.g., opening a window in VR).
@@ 4.2. Case 1: How 911 changed Financial Distract in Manhattan
This story covers the societal changes in Manhattan’s Financial District post-9/11 (911). The original story featured a 3D model of Manhattan’s densely packed buildings, making it ideal for an exocentric VR adaptation. The story has 15 steps and 611 words in total. A step includes a paragraph of story narratives with corresponding visual elements. Audiences procceed to next step by following a specific navigation trigger.
Main Visual Element. We built a 3D landscape of Manhattan in VR (Figure 5-N1) through Mapbox (mapbox_2017_mapbox). The landscape is 2m-by-2m square base map by default. The current building information came from mapbox.3d-buildings tileset. When the story narratives mentioned a particular group of buildings (e.g. WTC buildings), they are highlighted in blue.
@@ 4.3. Case 2: In the Atlantic Ocean, Subtle Shifts Hint at Dramatic Dangers
This story explores the impact of climate change on the Gulf Stream (atlantic). We only selected the first half of the story, which has the most concentrated 3D spatial information. The main visualization of the original story was a globe with animated current circulations in Atlantic Ocean. Thus, we implemented a corresponding exocentric VR story. We summarized the text-only portion (484 words) into 2 paragraphs (100 words) to keep the length consistency. The story has 15 steps and 503 words in total.
@@ 4.4. Case 3: What the Tulsa Race Massacre Destroyed
The story of this case illustrates the business achievement of black Americans in Tulsa, Oklahoma in the early 1900s and how the town was destroyed by racist mobs (tulsa). We only selected the first half the story, which has the most concentrated 3D information. We summarized the second half of the story into a paragraph (130 words) and put it at the end so that story narratives are logically complete while keeping similar length with previous cases. The main visualization of the original story comprises a 3D reconstruction of Tulsa before the massacre, which leads to our design of a VR story with an egocentric viewpoint in a main avenue (Greenwood Ave) of Tulsa (Figure 7-T3). The story has 18 steps and 618 words in total.
@@ 5.5. Task Data Collection
Through the study, we aimed to investigate the impact of ego/exocentric viewpoints and active/passive navigation on spatial immersion and understanding of VR stories. For spatial immersion, we collected responses to the Presence Questionnaire (berkman2021group) and NASA-TLX (hart1988development) to measure spatial presence and perceived workload, respectively.
@@ 5.6. Hypotheses
Story Session Duration. We hypothesize that active navigation will result in longer story session durations compared to passive navigation. The interactive elements in active navigation require users to engage physically with the environment, adding time to the experience. Prior studies have shown that increased interaction in VR generally extends session time due to higher engagement levels and the need for precise actions in immersive environments (lages2018move; servotte2020vr). This additional engagement likely leads participants to spend more time exploring the environment and interacting with the content, especially in active conditions.
Additionally, Ego viewpoints are likely to further increase session duration, as spatial navigation in egocentric environments tends to be more complex and challenging, requiring users to explore the scene more actively, which can lead to extended engagement (van2018virtual).
Story Understanding. We hypothesize that active navigation will outperform passive navigation on both content comprehension and spatial information depiction, as embodied interaction has been shown to improve cognitive engagement and memory retention (wilson2002six; makransky2019adding). Active navigation encourages physical movement, which enhances attention to both content and spatial information, leading to better comprehension and recall (servotte2020vr).
We also expect that exocentric viewpoints will result in better content comprehension because they provide an overview of the entire scene, facilitating the integration of key story elements and improving factual recall (yang2018maps). Egocentric viewpoints will lead to better spatial understanding, as being immersed in the scene helps participants build stronger spatial awareness and memory, which is supported by the ”memory palace” technique (legge2012building; krokos2019virtual).
Spatial Immersion. We expect that Ego viewpoints combined with active navigation will result in the highest levels of spatial immersion. Active navigation increases the sense of presence and control, while Ego viewpoints create a fully immersive experience by surrounding participants with 3D elements (slater2009place). Prior studies suggest that immersion is enhanced when users feel directly involved in the environment, with Ego viewpoints offering a more embodied experience (cummings2016immersive).
While active navigation may increase perceived workload due to the effort required to interact with the environment, we hypothesize that the enhanced sense of presence and engagement outweighs the additional cognitive demand, especially when paired with the Ego viewpoint (hart1988development). This combination offers an optimal balance between immersion and interactivity, despite a potential trade-off with workload (servotte2020vr).
@@ 7.1. How does viewpoints influence story memorability?
Our study shows that the viewpoint is a more decisive factor than the navigation in story memorability. The exocentric viewpoint is slightly better than the egocentric one in the overview and detail memorization. This result matches the disccovery of some prior works, in which the exocentric viewpoint benefits analytical tasks in immersive environments (yang2018maps; kraus2019impact; yang2020embodied), but the advantage in our study is not as significant. Our anticipated reason is that the exocentric viewpoint provide comprehensive information within audiences’ view, which alleviates the problem of targeting story elements in space. However, this also causes a larger information density and more occlusions than the egocentric viewpoint. The relatively smaller scale of visual representations in exo viewpoints increases the attention to the story textual content received and prolongs the retention time in the memory so that they can did a better content comprehension. However, at the same time, audiences could become less sensitive to content transitions and therefore they might miss some important details. Overall, our results indicate that the exocentric viewpoint can be a better design if the primary goal is to promote the understanding and retention of narratives, and aforementioned problems should be considered to further improve the story.
In contrast, the egocentric viewpoint demonstrated better memorability on spatial information than exocentric viewpoint. Similar result is also found in Krokos et al.’s (krokos2019virtual) and Yang et al.’s (yang2020virtual) work. Objects in an egocentric viewpoin are often larger and more visible than the exocentric one. Resultedly, audiences can better recognize transforms (i.e. position, rotation and scale) and capture their changes. More importantly, audiences are involved as part of a egocentric story. This enable them to use themselves as a frame of reference to infer transforms. In exocentric stories, they have to refer to one central element (e.g. the base map in subsection 4.2) whose signficant transform change can greatly distract the spatial information perception since audiences need to find a new frame of reference every time. Thus, the egocentric viewpoint is beneficial for those authors who want to convey understandings and perceptions related to the changes 3D spatial transforms (e.g. the renovation process of an office).
@@ 7.2. How does navigation influence spatial immersion?
This finding aligns with Lages et al.’s study (lages2018move), which showed that walking-based navigation improved performance for users who lacked proficiency in VR manipulation. Our results suggest that the need for higher precision in finding specific angles in exocentric scenarios (e.g., subsection 4.2, subsection 4.3, and the grab-and-move tasks in Lages et al.’s study) may explain the increased workload. The “changing frame of reference” issue discussed in subsection 7.1 adds to the demand for spatial awareness and motor skills to achieve the correct angle, resulting in higher effort and frustration, as reflected in the study.
@@
When designing navigation for VR stories, authors should consider the story’s viewpoint based on these findings. However, focusing solely on individual navigation steps is insufficient. Ferguson et al. (ferguson2020role), in their investigation of VR educational games, recommend that active navigation should seamlessly integrate with the story content and involve minimal learning curves. We argue that natural active navigation in VR storytelling should combine story progression with user exploration.
@@ 8. Limitations and Future Work
We adapted four selected web-based stories into immersive environments. To guide this adaptation, we conducted a formative user study to identify critical design considerations and explored various design options. Our exploration was not exhaustive; it aimed to facilitate the adaptation process by focusing on standard, off-the-shelf VR solutions. This project concentrated on specific aspects of viewpoint and navigation: ego vs. exo and active vs. passive. While we believe these are unique and essential areas to investigate, we acknowledge a much broader design space for immersive storytelling beyond these aspects. For instance, more sophisticated VR locomotion techniques (di2021locomotion), such as redirected walking, could enhance the active navigation experience. Similarly, advanced visual guidance methods (lange2020hivefive), such as those for out-of-view scenarios (petford2019comparison; lin2021labeling) and multisensory approaches (melo2020multisensory), offer additional potential for enhancing user experience. We did not incorporate these techniques, as our primary focus was the user study and providing easily accessible techniques available on standard platforms. Building a comprehensive design space to inform future immersive story design is crucial, and we consider this an area for future work.
@@
Our study materials were adapted from web-based stories that were initially designed for web browsers and contained design patterns specifically tailored for web environments (bach2018narrative). It is unclear whether these design patterns are still applicable in immersive environments or whether we need to develop unique and more appropriate design patterns for immersive stories. Studying the adaptation or innovation of these design patterns for immersive stories has the potential to bring about a more native immersive experience. Furthermore, the web-based stories we adapted primarily follow a linear, single-scene structure. However, complex stories with multi-path or branching narratives also exist. While our studied scenario can serve as a foundational unit for these more complex stories, and our findings should be partially applicable, especially within their individual scenes. There is a need to explore transitions between scenes and more intricate active navigation, such as manipulating and selecting narrative branches.
@@
Through our study, we found that exocentric and egocentric viewpoints each have their own strengths and limitations. We considered viewpoint as an exclusive design element, testing it in a controlled manner. There is a great opportunity to design immersive stories that leverage both viewpoints to complement each other. In movies and data videos, cinematic techniques like camera movements and zooms have been used to switch focuses and direct audiences’ attention (amini2015understanding; xu2022from; li2023geocamera). Similar incoporation such as viewing some story pieces from an exocentric view and others from an egocentric view is also promising to investigate. Our study results can provide guidelines on which parts of the story to use which viewpoint. To achieve this, smooth transition techniques are necessary to allow seamless and meaningful switching between viewpoints, which we aim to explore in our next research phase.
@@
Finally, one important aspect of active vs. passive interaction is the agency of the readers. In our study, readers were given the freedom to navigate by themselves with guidance. However, there are other types of freedom that could be provided to readers, such as allowing them to freely decide their focus or even rearrange the storyline. This may empower readers but also risks them getting lost, particularly due to the more flexible interactivity in immersive environments. Balancing exploration and narrative control is a fundamental question in storytelling (thudt2018exploration). Studying the appropriate threshold or mechanism for this balance is an important research thread in immersive storytelling.